The electrical power for an implanted medical device decreases by 0.1% each day.

a. Find a formula for the nth term of the geometric sequence that gives the percent of the initial

power n days after the device is implanted.


1
Expert's answer
2021-11-15T16:41:23-0500

"1-0.1\% = 1-0.001 =0.999"

After "n" days the value of the device depreciates "n" times with a factor 0.999. The value of the device after "n" days is


"a_n=(initial value) (common ratio)^n"

"=1(0.999)^n = (0.999)^n"

The values of the device form a geometric sequence with "a_1=0.999" and "r =0.999"

The formula for the "n" th term of the geometric sequence that gives the percent of the initial power "n" day after the device is implanted is "a_n=(0.999)^n"